OpenStudy (anonymous):

evaluate the integrals:>> integral sign sin 3x dx

hartnn (hartnn):

do you know \[\int \sin ydy =... ?\]

OpenStudy (aravindg):

put t=3x

OpenStudy (aravindg):

then just do normal integration

OpenStudy (anonymous):

ahm..."???"

hartnn (hartnn):

yes, first tell what u get after u put t=3x dt =... ?

OpenStudy (anonymous):

3dx

hartnn (hartnn):

yes, so dx = dt/3 so whats your new integral ?>

hartnn (hartnn):

\[(1/3) \int \sin t dt = .. ?\]